# Modelling reliability of reversible circuits with 2D second-order cellular automata

### 20 Dec 2023 v1

@@ Abstract
The cellular automata is a widely known model of reversible computations. The family of reversible second-order cellular automata considered in this work is appropriate both for construction of logic gates and analysis of damage distribution. The quantities such as formal dimension of damage patterns can be used only for rough estimation of consequences of particular faults and numerical experiments are provided for illustration of some subtleties. Such analysis demonstrates high sensitivity to errors from defects, lack of synchronization and too short intervals between signals.
Graphical abstract [Uncaptioned image]
@@ 1 Introduction
The reversible computations have quite a long history [1] and nowadays they deserve increasing attention due to different applications from the reduction of power consumption to the quantum computations [2, 3, 4, 5]. The analysis of uncertainties for the testing of usual logic circuits is well developed area with variety of different methods [6], but some subtleties can appear for reversible circuits. The problems of faults and the reliability of the reversible circuits were also covered quite widely [7, 8, 9].
@@
From the one side the testing of reversible circuits may be even simpler for some particular faults [7], but some difficulties may appear in the more general cases [8]. There are also some problems due to specific properties of reversible circuits. From the one hand, the reversibility produces an advantage due to negligible power consumption, from the other one, the undesirable perturbations may spread without dissipation and produce the problems with too high error sensitivity.
@@
More detailed analysis of the particular structures used for construction of the reversible logical circuits can be useful for the sensitivity analysis and consideration of some other problems. A model of so-called quantum-dot cellular automata can be used as some example [9, 10]. The presented paper is devoted to the more abstract models based on the reversible cellular automata (RCA).
@@
The consideration of the uncertainty propagation can be described with irreversible probabilistic cellular automata [11], however an application of similar ideas to the reversible case produces some problems if to try add the noise directly into update rule of the cells, because such a system could not perform the reliable computations [12].
@@
The issue with noisy local rules of the probabilistic RCA justifies sensitivity analysis of usual (deterministic) RCA evolution with respect to different perturbations of the initial configurations. The family of second-order RCA with four states considered in this work uses the standard method of construction from the cellular automata with two states [13, 14, 15].
@@
The presented work is in some agreement with an earlier suggestion from Ref. [15] ‘[…] the most practical approach is to start with a local map that directly supports logic gates and wires, and then build the appropriate logic circuits out of these primitives’. Examples of such approach often inspired by so-called billiard ball model (BBM) [16]. A computationally universal second-order RCA derived from cellular automaton with three states was discussed already in Ref. [13].
@@
Another approach to the computational universality in RCA uses 1D design based on universal Turing machine instead of logic circuits [2, 17]. However, the model with 2D RCA and logic circuits discussed below provides a natural way to estimate sensitivity and reliability of reversible circuits.
@@
Let us briefly outline contents of this paper. A family with a few similar RCA and constructions of few logic gates are introduced in Sec. 2. Different kinds of errors from the defects and the improper signal interactions are discussed in Sec. 3 recollecting some quantities such as Lyapunov exponents and formal dimensions those could be useful for analysis of the damage spread together with the numerical experiments. In conclusion, Sec. 4 summarizes this work.
@@ 2.1 Basic properties and classification of CA
Let us recollect some properties of cellular automata (CA) on the rectangular grids with dimension $d$ denoted further as $\mathbb{Z}^{d}$ [20, 21]. Two-dimensional CA are quite natural for modelling of the logic circuits. The state of any cell on such a grid belongs to the finite-dimensional set $S$ with $k$ elements and for simplicity may be described by a number from interval $10,\ldots,k-$ 1. State $k=0$ could be considered as an empty cell and there are $k-1$ kinds of occupied cells. In the simplest case with two states the cell is either occupied or not.
@@
The reversible cellular automata (RCA) are appropriate for the purposes of presented work, but there are no effective algorithms for inverting given local rule for $d>1$ [17]. However, some methods allow to construct new local rules with known simple inverse from the very beginning.
One of these methods uses so-called second-order CA [13, 14, 15, 17, 18] recollected below. The second-order CA uses values from two previous steps for calculating of a state for a next one
@@
where $\mathbin{\ominus_{k}}$ is subtraction modulo $k$ and such CA is reversible [13, 14, 15, 18] with the inverse map
@@
For an initial CA with two states $k=2$ the operation $\mathbin{\ominus_{2}}$ coincides with XOR (eXclusive OR often denoted as $\oplus$) producing more familiar examples [17]. The second-order RCA can be considered as first-order RCA with state spaces $S\times S$ and local rule [15, 17]
@@ 2.2 Classification of CA and RCA
Due to Eq. (5) there is one to one correspondence between the second-order RCA with four states and two-state CA. Thus, classification of 2D CA with two states can be also used for such RCA. A natural choice is the isotropic local rules also known as completely symmetric [22] i.e., invariant with respect to rotations and reflections. However, even with such restriction there are $2^{102}\simeq 5\times 10^{30}$ different CA [22].
There are more narrow classes of CA, for example totalistic with value $f$ in Eq. (1) is depending only on sum of occupied cells in neighborhood and outer totalistic with taking into account the sum and the value of cell itself [22]. The last example is also known as semi-totalistic or Life-like CA [23, Ch. 6] due to famous Conway’s Game of Life CA [24, 25] and there are $2^{18}\simeq 3\times 10^{5}$ such rules [22].
For the certainty the term outer totalistic is saved further for particular case of outer totalistic inner-independent CA [23, Ch. 13] with local rule taking into account only ‘outer neighborhood’ without cell itself.
Between isotropic and totalistic classes there is an intermediate one with separate counts for two kinds of neighboring cells: the four closest neighbors have a common edge with central cell and another four cells have only common corner. Such a class is also known as corner-edge totalistic (CET) or quarter-totalistic CA rules [26] and it is convenient again to consider inner-independent case without dependence on central cell denoted further as corner-edge outer totalistic (CEOT) CA.
@@
The similar method can be used to calculate number of possible local rules in other classes of CA with two states discussed above and corresponding second-order RCA. It should be mentioned, that the local rule of RCA derived from CEOT CA due to Eq. (5) anyway uses value of central cell, but consideration only CEOT CA still can be reasonable, because number of RCA derived from CET CA is much bigger and more difficult for investigation.
The considered classes also can be convenient due to relations with invariant quantities in second-order RCA defined by count of cell values [14, 18, 27, 28, 29]. However, there are few alternative methods of classifications [30] and these topics should be discussed elsewhere.
@@ 2.3 RCA family for circuits construction
Let us consider family of CEOT CA with local update rules producing occupied cell only if all cells with common corner are empty, i.e., only pairs with $n_{c}=0$ may appear in the sequence for $f(p)=1$ discussed above. The simple example of such CA denoted here as $C_{0}E_{1}$ includes single pair with $n_{c}=0$, $n_{e}=1$ and second-order RCA derived from such a rule using Eq. (5) is denoted as $C_{0}E_{1}^{\,I\!I}$.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ also has moving configurations appropriate for encoding of signals in the circuits. There are infinite amount of possible signals represented as lines with length $k\leq 3$ starting with red cell (state $s=1$) and ending with a green cell (state $s=2$) with $k-2$ blue cells (state $s=3$) in between (‘snake’). If to use second-order description with two layers such configurations correspond to two lines with length $k-1$ with one cell displacement in direction of motion. The speed of such signals is one cell in vertical or horizontal direction per time step.
@@ 2.4 Constructions of reversible gates
An analogy of BBM computations [16] can be developed with shortest three-cell signals (‘ants’). The simplest encoding is supposed in such a case with the unit and zero are represented by presence or absence of an elementary signal respectively. It has some disadvantages, but demonstrates many basic principles of design and more difficult encodings should be discussed elsewhere.
@@
The important element for BBM is the reversible Fredkin gate [16]. Such a gate (with three inputs and three outputs) performs conditional exchange between two data signals depending on value of a control one. The action of Fredkin gate can be formally expressed as
@@
Here two data signals are exchanged if control signal presents. Such convention formally differs from initially suggested in Ref. [16], but it can be simply altered by exchange of data outputs.
The exchange (swap) gate is a simplest element, but for 2D CA some efforts are required to prevent collision of signals during exchange. Two possible designs with ‘wide’ and ‘narrow’ swap gates are shown on Fig. 1.
Figure 1: Swap gates: ‘wide’ (left) and ‘narrow’ (right)
@@
The pictures Fig. 1 also illustrate specific way of interaction between signals and static elements: there is always an one-cell gap between them. The signal stops, because new cell may not appear near static elements and starts motion in an orthogonal direction. To prevent unlimited growth one of the directions should be limited by a design of the static elements.
The Fredkin gates together with delay elements, swaps and ‘constants’ (i.e. some amounts of auxiliary signals) allow to construct any logic circuits [16]. Necessity of some constants for used encoding is clear even from construction of NOT gate that can be expressed with Fredkin gate Eq. (6) as
@@
Similarly with BBM construction of the Fredkin gate can be expressed with simpler switch gates [16] for conditional routing of one data signal by a control one. The scheme of switch gate configuration for discussed RCA is shown on Fig. 2.
Figure 2: Signals routing and collisions in a model of switch gate
@@
The scheme of Fredkin gate composed from four such switch gates initially proposed by R. Feynman and A. Ressler may be found in Ref. [16]. Possible realization for considered RCA is shown on Fig. 3. Without control signal two data signals follow pair of paths and simply leave the scheme with some delay, but due to interaction with control signal data signals follow other paths to exchange output positions. Many static element here is for synchronization of delays for arbitrary combinations of input signals.
Figure 3: Signals routing and collisions in a model of Fredkin gate
@@
Such property provides a counterexample to some confusion about impossibility to split signals in reversible system. The inversion of splitting element is merging of two signals. Let us consider yet another element useful for construction of Controlled-NOT gate. Similarly with switch gate it has two inputs and three output paths, but it may merge two of signals into one path. It encodes three possible non-empty combinations of two signals into one signal moving along one of three possible paths, see Fig. 4.
@@
If to consider Controlled-NOT gate as a function on pairs of bits, it exchanges rows $(1,0)$ and $(1,1)$ in the table above. It can be implemented by encoding pairs of signals into three different paths using scheme on Fig. 4, exchanging two appropriate paths and applying the inverse of the scheme (with split of signal along one of paths). Possible design is shown on Fig. 5.
@@
Thus, considered family of CEOT RCA defined by property $n_{c}=0$ can be divided on two types with similar regular behavior, but with some subtleties relevant for analysis of sensitivity and reliability. The first type may use fixed number of signals with different length and for the second one the length is fixed, but signals can be split and merged.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ was briefly mentioned first in Ref. [31], $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ were introduced and implemented in software developed by author with different names around the same time [32] together with yet another isotropic RCA.
@@
It was already mentioned, that the second-order RCA can be considered as usual (first-order) RCA with four states. Using conversion into appropriate format for local rules with four states some numerical experiments discussed here was also performed on general software for CA simulation developed by A. Trevorrow, T. Rokicki et al [33].

### 19 Mar 2024 v2

@@ Abstract
The cellular automaton is a widely known model of both reversible and irreversible computations. The family of reversible second-order cellular automata considered in this work is appropriate both for construction of logic gates and analysis of damage distribution. The quantities such as formal dimension of damage patterns can be used only for rough estimation of consequences of particular faults and numerical experiments are provided for illustration of some subtleties. Such analysis demonstrates high sensitivity to errors from defects, lack of synchronization and too short intervals between signals.
keywords:
reversible computations, cellular automata, damage spread
journal: arXiv{graphicalabstract}
@@ 1 Introduction
Reversible computations have quite a long history Ben73 and nowadays they deserve increasing attention due to different applications from the reduction of power consumption to the theory of quantum computations KM17 ; COST20 ; RC22 ; RC23 . The analysis of uncertainties for the testing of usual logic circuits is well developed area with variety of different methods Circ13 , but some subtleties can appear for reversible circuits. The problems of faults and the reliability of the reversible circuits were also covered quite widely PHM03 ; HPB04 ; RTest20 .
@@
From the one side the testing of reversible circuits may be even simpler for some particular faults PHM03 , but some difficulties may appear in the more general cases HPB04 . There are also some problems due to specific properties of reversible circuits. From the one hand, the reversibility produces an advantage due to negligible power consumption, from the other one, the undesirable perturbations may spread without dissipation and produce the problems with too high error sensitivity.
@@
More detailed analysis of the particular structures used for construction of the reversible logical circuits can be useful for the sensitivity analysis and consideration of some other problems. A model of so-called quantum-dot cellular automata can be used as some example RTest20 ; QDCA07 . The presented paper is devoted to the more abstract models based on the reversible cellular automata (RCA).
@@
The consideration of the uncertainty propagation can be described with irreversible probabilistic cellular automata UQCA , however an application of similar ideas to the reversible case produces some problems if to try add the noise directly into update rule of the cells, because such a system could not perform the reliable computations RCAnoise .
@@
The issue with noisy local rules of the probabilistic RCA justifies sensitivity analysis of usual (deterministic) RCA evolution with respect to different perturbations of the initial configurations. The family of second-order RCA with four states considered in this work uses the standard method of construction from the cellular automata with two states M84 ; V84 ; TM90 .
@@
The presented work is in some agreement with an earlier suggestion from Ref. TM90 ‘[…] the most practical approach is to start with a local map that directly supports logic gates and wires, and then build the appropriate logic circuits out of these primitives’. Examples of such approach often inspired by so-called billiard ball model (BBM) KM17 ; FT82 ; JK18 . A computationally universal second-order RCA derived from cellular automaton with three states was discussed already in Ref. M84 .
@@
Another approach to the computational universality in RCA uses 1D design based on universal Turing machine and it is discussed together with 2D logic circuits based on so-called partitioned CA (PCA) in Ref. KM17 ; JK18 . PCA is usual choice for modelling of reversible circuits due to natural relation with BBM.
The conservation of number of occupied cells can be simply implemented in PCA directly by local transition rules preventing some problems with damage distribution discussed further. On the other hand, more realistic models of logic circuits may require possibility to investigate noise and damage.
Models without conservation of cell number may be necessary in such a case. In irreversible CA undesired fast growth of number of cells could be corrected by modification of local rules to cancel some configurations. For example, diverse evolution Game of Life CA WW ; Gard and possibility of universal computations with such a model supported by rules with good balance between appearance and cancellation of different configurations.
However, there is specific problem with RCA, because reversible evolution make impossible extinction of any configuration and without conservation of cell number (or some other non-negative quantity) undesired growth and chaotic behaviour of initially regular configuration is rather common result of RCA evolution in unbounded domain.
Few reversible second-order RCA discovered by author some time ago and discussed in this work provide interesting combination of stable and expansive behaviour. From the one hand such RCA can be used for creation of reversible universal gates, from the other one even expansive behaviour displays some regularities and fractal patterns. Investigation sensitivity and reliability of reversible circuits with such models can be quite instructive and relevant to analogue processes in more realistic systems.
@@
Let us briefly outline contents of this paper. After recollection of basic properties and possible classification of CA in Sec. 2.1 and Sec. 2.2 few original second-order RCA for circuits construction are introduced in Sec. 2.3 and examples of logic gates for universal computations are described in Sec. 2.4. Different kinds of errors from the defects and the improper signal interactions are discussed in Sec. 3 recollecting some quantities such as Lyapunov exponents and formal dimensions those could be useful for analysis of the damage spread together with the numerical experiments. In conclusion, Sec. 4 summarizes this work.
@@ 2.1 Basic properties and classification of CA
Let us recollect some properties of cellular automata (CA) on the rectangular grids with dimension $d$ often denoted as $\mathbb{Z}^{d}$ WolCA ; JK05 . Two-dimensional CA are quite natural for modelling of the logic circuits. The state of any cell on the grid $\mathbb{Z}^{d}$ belongs to the finite-dimensional set $S$ with $k$ elements and for simplicity may be described by a number from interval $10,\ldots,k-$ 1. State $k=0$ could be considered as an empty cell and there are $k-1$ kinds of occupied cells. In the simplest case with two states the cell is either occupied or not.
@@
The reversible cellular automata (RCA) are appropriate for the purposes of presented work, but there are no effective algorithms for inverting given local rule for $d>1$ JK18 . However, some methods allow to construct new local rules with known simple inverse from the very beginning.
One of these methods uses so-called second-order CA M84 ; V84 ; TM90 ; JK18 ; Ila recollected below. The second-order CA uses values from two previous steps for calculating of a state for a next one
@@
where $\mathbin{\ominus_{k}}$ is subtraction modulo $k$ and such CA is reversible M84 ; V84 ; TM90 ; Ila with the inverse map
@@
For an initial CA with two states $k=2$ the operation $\mathbin{\ominus_{2}}$ coincides with XOR (eXclusive OR often denoted as $\oplus$) producing more familiar examples JK18 . The second-order RCA can be considered as first-order RCA with state spaces $S\times S$ and local rule TM90 ; JK18
@@ 2.2 Classification of CA and RCA
Due to Eq. (5) there is one to one correspondence between the second-order RCA with four states and two-state CA. Thus, classification of 2D CA with two states can be also used for such RCA. A natural choice is the isotropic local rules also known as completely symmetric PW85 i.e., invariant with respect to rotations and reflections. However, even with such restriction there are $2^{102}\simeq 5\times 10^{30}$ different CA PW85 .
There are more narrow classes of CA, for example totalistic with value $f$ in Eq. (1) is depending only on sum of occupied cells in neighborhood and outer totalistic with taking into account the sum and the value of cell itself PW85 . The last example is also known as semi-totalistic or Life-like CA (GoL10, , Ch. 6) due to famous Conway’s Game of Life CA WW ; Gard and there are $2^{18}\simeq 3\times 10^{5}$ such rules PW85 .
For the certainty the term outer totalistic is saved further for particular case of outer totalistic inner-independent CA (GoL10, , Ch. 13) with local rule taking into account only ‘outer neighborhood’ without cell itself.
Between isotropic and totalistic classes there is an intermediate one with separate counts for two kinds of neighboring cells: the four closest neighbors have a common edge with central cell and another four cells have only common corner. Such a class is also known as corner-edge totalistic (CET) or quarter-totalistic CA rules term and it is convenient again to consider inner-independent case without dependence on central cell denoted further as corner-edge outer totalistic (CEOT) CA.
@@
The similar method can be used to calculate number of possible local rules in other classes of CA with two states discussed above and corresponding second-order RCA. The considered classes also can be convenient due to relations with invariant quantities in second-order RCA defined by count of cell values V84 ; Ila ; YP84 ; GV86 ; ST12 . However, there are few alternative methods of classifications class and these topics should be discussed elsewhere.
@@ 2.3 RCA family for circuits construction
Let us consider subfamily of CEOT CA with local update rules producing occupied cell only if all cells with common corner are empty, i.e., only pairs with $n_{c}=0$ may appear in the sequence for $f(p)=1$ discussed above. Selected CA are between simplest examples of CA from considered families (CET, CEOT) and were found by not very extensive search for CA with universal gates. Anyway, analogue CA were not found in similar way between totalistic rules and so CEOT was used instead. The simple example of such CA denoted here as $C_{0}E_{1}$ includes single pair with $n_{c}=0$, $n_{e}=1$ and second-order RCA derived from such a rule using Eq. (5) is denoted as $C_{0}E_{1}^{\,I\!I}$. Few different extensions of such RCA defined further could be denoted as $C_{0}E_{1\dots}^{\,I\!I}$.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ also has moving configurations appropriate for encoding of signals in the circuits. There are infinite amount of possible signals represented as lines with length $k\geq 3$ starting with red cell (state $s=1$) and ending with a green cell (state $s=2$) with $k-2$ blue cells (state $s=3$) in between (‘snake’). If to use second-order description with two layers such configurations correspond to two lines with length $k-1$ with one cell displacement in direction of motion. The speed of such signals is one cell in vertical or horizontal direction per time step.
@@ 2.4 Constructions of reversible gates
An analogy of BBM computations FT82 can be developed with shortest three-cell signals (‘ants’). The simplest encoding is supposed in such a case with the unit and zero are represented by presence or absence of an elementary signal respectively. It has some disadvantages, but demonstrates many basic principles of circuits design and more difficult encodings should be discussed elsewhere.
@@
The important element for BBM is the reversible Fredkin gate FT82 . Such a gate (with three inputs and three outputs) performs conditional exchange between two data signals depending on value of a control one. The action of Fredkin gate can be formally expressed as
@@
Here two data signals are exchanged if control signal presents. Such convention formally differs from initially suggested in Ref. FT82 , but it can be simply altered by exchange of data outputs.
The exchange (swap) gate is a simplest element, but for 2D CA some efforts are required to prevent collision of signals during exchange. Two possible designs with ‘wide’ and ‘narrow’ swap gates are shown on Figure 1.
Figure 1: Swap gates for $C_{0}E_{1\dots}^{\,I\!I}$: ‘wide’ (left) and ‘narrow’ (right)
@@
The pictures Figure 1 also illustrate specific way of interaction between signals and static elements: there is always an one-cell gap between them. The signal stops, because new cell may not appear near static elements and starts motion in an orthogonal direction. To prevent unlimited growth one of the directions should be limited by a design of the static elements.
The Fredkin gates together with delay elements, swaps and ‘constants’ (i.e. some amounts of auxiliary signals) allow to construct any logic circuits FT82 . Necessity of some constants for used encoding is clear even from construction of NOT gate that can be expressed with Fredkin gate Eq. (6) as
@@
Similarly with BBM construction of the Fredkin gate can be expressed with simpler switch gates FT82 for conditional routing of one data signal by a control one. The configuration of the switch gate for discussed RCA is shown on the left side of Figure 2. The scheme of the switch gate shown on the right side of the figure has been modified compared to that used in FT82 to more accurately display the directions of outgoing signals.
Figure 2: Signals routing and collisions in a model of switch gate for $C_{0}E_{1\dots}^{\,I\!I}$
@@
Figure 3: Signals routing and collisions in a model of Fredkin gate for $C_{0}E_{1\dots}^{\,I\!I}$
The scheme of Fredkin gate composed from four such switch gates initially proposed by R. Feynman and A. Ressler may be found in Ref. FT82 . Possible realization for considered RCA is shown on Figure 3. The simplified circuit diagram without delay elements shown at the bottom of Figure 3 uses two switch gates together with two inverted switch gates. All switch gates on Figure 3 are flipped vertically compared with Figure 2 and surrounded by dashed rectangles on the diagram.
Without control signal two data signals follow pair of bypass paths (above the main diagram) and simply leave the scheme with some delay, but due to interaction with control signal data signals follow other (triple-crossing) paths to exchange output positions. Many static element here is for synchronization of delays for arbitrary combinations of input signals.
@@
Such property provides a counterexample to some confusion about impossibility to split signals in reversible system. The inversion of splitting element is merging of two signals. Let us consider yet another element useful for construction of Controlled-NOT gate. Similarly with switch gate it has two inputs and three output paths, but it may merge two of signals into one path. It encodes three possible non-empty combinations of two signals into one signal moving along one of three possible paths, see Figure 4. The simplified scheme of such gate is shown on the right side of the figure.
@@
If to consider Controlled-NOT gate as a function on pairs of bits, it exchanges rows $(1,0)$ and $(1,1)$ in the table above. It can be implemented by encoding pairs of signals into three different paths using scheme on Figure 4, exchanging two appropriate paths and applying the inverse of the scheme (with split of signal along one of paths).
Possible design is shown on Figure 5. The simplified circuit diagram without delay elements shown at the bottom of Figure 5 uses a gate depicted on Figure 4 (surrounded by a dashed rectangle) together with inversion of such gate and exchanges paths of two signals between them.
@@
The Toffoli gate also can be designed for such RCA as $C_{0}E_{1,2}^{\,I\!I}$, but elaborated configuration is not shown here due to typographical and other reasons. Anyway, the Toffoli gate can be composed from Fredkin and Controlled-NOT gates shown earlier on Figure 3 and Figure 5.
Thus, considered family of CEOT RCA defined by property $n_{c}=0$ can be divided on two types with similar regular stable behaviour, but with some subtleties with expansive behaviour discussed further and relevant for analysis of sensitivity and reliability. The first type may use fixed number of signals with different length and for the second one the length is fixed, but signals can be split and merged.
@@
The RCA $C_{0}E_{1}^{\,I\!I}$ was briefly mentioned first in Ref. AV14 , $C_{0}E_{1,2}^{\,I\!I}$, $C_{0}E_{+}^{\,I\!I}$ and $C_{0}E_{1-3}^{\,I\!I}$ were introduced and implemented in software developed by author with different names around the same time CART together with yet another isotropic RCA.
@@
It was already mentioned, that the second-order RCA can be considered as usual (first-order) RCA with four states. Using conversion into appropriate format for local rules with four states some numerical experiments discussed here was also performed on general software for CA simulation developed by A. Trevorrow, T. Rokicki et al GoL .